OpenStudy (anonymous):

Suppose f(x) is continuous on [2,8] and −4≤f′(x)≤5 for all x in (2,8). Use the Mean Value Theorem to estimate f(8)−f(2).

OpenStudy (experimentx):

could be between -24 to 60

OpenStudy (anonymous):

-24 works

OpenStudy (experimentx):

??

OpenStudy (anonymous):

-24 to 30

OpenStudy (anonymous):

experimentx's typo :)

OpenStudy (experimentx):

any idea chlorophyll??

OpenStudy (anonymous):

thanks chlorophyll

OpenStudy (anonymous):

one last question By applying Rolle's theorem, check whether it is possible that the function f(x)=x^9+x−18 has two real roots. Answer: (input possible or impossible )...ans is impossible Your reason is that if f(x) has two real roots then by Rolle's theorem: f′(x) must be (input a number here)

OpenStudy (anonymous):

-4 = < Δy/ Δx <= 5 -4 * 6 = < Δy < = 5 * 6

OpenStudy (anonymous):

i need help wit the second part pls

OpenStudy (experimentx):

you sure the answer is that??

OpenStudy (anonymous):

-24 t0 30 worked

myininaya (myininaya):

\[2 \le x \le 8\] \[-4 \le f'(x) \le 5\] f is continuous in differentable on [2,8] so we know there is c btw 2 and 8 such that \[f'(c)=\frac{f(8)-f(2)}{8-2}\] And we know \[-4 \le f'(c) \le 5 \] => we have \[-4 \le \frac{f(8)-f(2)}{8-2} \le 5 \]

myininaya (myininaya):

\[-4 (8-2) \le f(8)-f(2) \le 5(8-2)\]

myininaya (myininaya):

and not in* sorry type-o there my english sucks but the math is good lol

OpenStudy (anonymous):

lol thanks. can you help me with this last one

OpenStudy (anonymous):

By applying Rolle's theorem, check whether it is possible that the function f(x)=x^9+x−18 has two real roots. Answer: (input possible or impossible )...ans is impossible Your reason is that if f(x) has two real roots then by Rolle's theorem: f′(x) must be (input a number here

OpenStudy (experimentx):

was it a parabola?? i thought i saw ^9 up there

OpenStudy (experimentx):

well, that was interestig, f(c) would give the value of vertex. if negative, and the parabolas is concave upwards then two real roots.

OpenStudy (experimentx):

and if f(c) positive and the paraboal concave downwards then also two real roots. if f(c) is zero than single roots ... but i was wondering where would we apply Rolle's theorem.

OpenStudy (experimentx):

i guess it would be just to assume that f(a) = f(b) = 0 , and a,b be the roots real roots i.e. point c implies point a and point b on x-axis?? but what to do if there're were single root??

OpenStudy (anonymous):

zero works thanks very much for the help

OpenStudy (experimentx):

point a, b, and c points coincide ... well that atleast work for parabola.

OpenStudy (experimentx):

the real problem lies in this type of graph  that was a parabola, we knew that if it crosses x-axis then it means either it opens upward or downward ... but i really don't know what to do about this type of graph.